# On Effects of Steering Latent Representation for Large Language Model Unlearning

### v1

@@ Abstract
Representation Misdirection for Unlearning (RMU), which steers model representation in the intermediate layer to a target random representation, is an effective method for large language model (LLM) unlearning. Despite its high performance, the underlying cause and explanation remain underexplored. In this paper, we first theoretically demonstrate that steering forget representations in the intermediate layer reduces token confidence, causing LLMs to generate wrong or nonsense responses. Second, we investigate how the coefficient influences the alignment of forget-sample representations with the random direction and hint at the optimal coefficient values for effective unlearning across different network layers. Third, we show that RMU unlearned models are robust against adversarial jailbreak attacks. Last, our empirical analysis shows that RMU is less effective when applied to the middle and later layers in LLMs. To resolve this drawback, we propose Adaptive RMU—a simple yet effective alternative method that makes unlearning effective with most layers. Extensive experiments demonstrate that Adaptive RMU significantly improves the unlearning performance compared to prior art while incurring no additional computational cost.
@@ 1 Introduction
State-of-the-art LLMs such as GPT-4 (Achiam et al. 2023), Gemini (Team et al. 2023), Llama-3 (Meta 2024), and Claude-3 Sonnet (Anthropic 2024) achieve remarkable performance through pre-training on large amounts of internet texts and rigorous alignment process for safety enhancement. Despite the immense effort in safety research, LLMs are still vulnerable to adversarial jailbreak attacks and can exhibit unwanted behaviors (Shah et al. 2023; Chao et al. 2023; Zou et al. 2023b; Jones et al. 2023; Yuan et al. 2024; Wei, Haghtalab, and Steinhardt 2024).
Machine Unlearning (Cao and Yang 2015; Chris Jay Hoofnagle and Borgesius 2019; Bourtoule et al. 2021; Nguyen et al. 2022; Xu et al. 2023; Liu et al. 2024c) has emerged as a promising method for mitigating unforeseen risks in LLMs before deployment. Li et al. (2024b) introduced Representation Misdirection for Unlearning (RMU)—an unlearning method that steers the representations of forget-samples (i.e. samples that the model should forget) toward a random representation while keeping the representations of retain-samples (i.e. samples that the model should remember) unchanged. RMU significantly degrades models’ accuracy on forget-tasks, while only slightly affecting the performance on retain-tasks and demonstrates stronger robustness against adversarial jailbreak attacks. However, the reason for RMU’s effectiveness is not well understood, hindering the development of better unlearning algorithms. In this paper, we make the following contributions:
@@
  • We investigate the connection between RMU and adversarial robustness. We demonstrate that RMU impedes the adversary’s ability to determine optimal updates for generating adversarial samples, thus improving the adversarial robustness of the model.
@@ 2 Background and related work
A natural is leave-some-out retraining: retraining the model from scratch without the forget samples. However, this method becomes more computationally expensive as the size of datasets and modern deep networks grows. Existing works focus on approximating unlearning (Warnecke et al. 2021; Izzo et al. 2021; Sekhari et al. 2021; Isonuma and Titov 2024) using Influence Function (Koh and Liang 2017; Grosse et al. 2023), gradient projection (Bae et al. 2023), gradient ascent (Thudi et al. 2022; Trippa et al. 2024), second-order approximation (Jia et al. 2024), preference optimization (Zhang et al. 2024b), and embedding corrupted (Liu et al. 2024a). Other views on the landscape of machine unlearning include: unlearning in text classification (Ma et al. 2022), image classification and recognition (Ginart et al. 2019; Golatkar, Achille, and Soatto 2020; Fan et al. 2024; Choi and Na 2023; Cha et al. 2024), image-to-image generative models (Li et al. 2024a), diffusion models (Gandikota et al. 2023; Zhang et al. 2024a; Kumari et al. 2023), multimodal unlearning (Cheng and Amiri 2023), federated unlearning (Liu et al. 2020a; Romandini et al. 2024; Wang et al. 2022; Che et al. 2023; Halimi et al. 2022; Jeong, Ma, and Houmansadr 2024), graph unlearning (Chen et al. 2022; Chien, Pan, and Milenkovic 2023; Wu et al. 2023a; Said et al. 2023; Cheng et al. 2023; Dukler et al. 2023; Zhu, Li, and Hu 2023; Li et al. 2024c; Tan et al. 2024), recommender systems (Zhang et al. 2023; Chen et al. 2024; Li et al. 2023; Wang et al. 2024), certified minimax unlearning (Liu et al. 2024b), and evaluation on unlearning (Lynch et al. 2024; Hayes et al. 2024; Shi et al. 2024a, b).
@@
Due to the large size of the parameters and training data, LLM poses a new challenge to unlearning. Current studies in LLM unlearning mainly focus on task or context-specific settings such as unlearning copyrighted material from the Harry Potter series (Eldan and Russinovich 2023), In-context unlearning (Pawelczyk, Neel, and Lakkaraju 2023), fictitious unlearning (Maini et al. 2024), specific harmful input-output (Yao, Xu, and Liu 2023; Liu et al. 2024d), sensitive and private information (Jang et al. 2023; Wu et al. 2023b; Ishibashi and Shimodaira 2023; Patil, Hase, and Bansal 2024), gender (Belrose et al. 2023) concepts (Hong et al. 2024), or facts (Meng et al. 2022). More recently, Li et al. (2024b) consider unlearning an entire distribution of hazardous knowledge given limited samples.
@@
Let $\mathcal{D}_{\text{forget}}$ and $\mathcal{D}_{\text{retain}}$ be the forget and retain sets, respectively. Let $f_{\theta}:\mathbb{R}^{n\times d}\mapsto\mathbb{R}^{n\times|V|}$ be an autoregressive LLM parameterized by $\theta$ that maps a prompt input $x_{1:n}$ consisting of $n$ tokens $\{x_{1},x_{2},...,x_{n}\}$ to an output of probability distributions over the vocabulary $V$. $h_{\theta}^{(l)}(x)$ denotes the averaged hidden states of input tokens $x_{i}$ from the $l$-th layer of $f_{\theta}$. Our goal is to unlearn the undesired harmful knowledge $\mathcal{D}_{\textnormal{forget}}$ from $f_{\theta}$ while retaining unrelated or general knowledge $\mathcal{D}_{\textnormal{retain}}$. Unlearned models should be robust to knowledge recovery attacks that attempt to recover harmful knowledge from the model.
Representation Misdirection for Unlearning.
(RMU; Li et al. (2024b)) is a fine-tuning-based unlearning method inspired by representation engineering (RepE; Zou et al. (2023a)) that steers the model’s representation of forget samples $x_{F}\in\mathcal{D}_{\text{forget}}$ to a random vector and regularizes the model representation of retain samples $x_{R}\in\mathcal{D}_{\text{retain}}$ back to the original model representation, by using the MSE loss:
$\mathcal{L}=||h_{\theta^{\text{unlearn}}}^{(l)}(x_{F})-c\bm{u}||_{2}^{2}+\alpha||h_{\theta^{\text{unlearn}}}^{(l)}(x_{R})-h_{\theta^{\text{frozen}}}^{(l)}(x_{R})||_{2}^{2}$ (1)
Where $\theta^{\text{unlearn}}$ and $\theta^{\text{frozen}}$ are parameters of the update model and frozen model respectively, $\bm{u}$ is a fixed random unit vector sampled from Uniform distribution $U(0,1)$, $c\in\mathbb{R}$ is a fixed scaling coefficient, and $\alpha$ is a retain weight. RMU updates $\theta^{\text{unlearn}}$ in the direction of the gradient of the loss $\mathcal{L}$ with respect to (w.r.t) $\theta$ using gradient descent.
@@ 3.1 The confidence of tokens generated by RMU models
In general, samples from the shifted distribution (such as wrong label or out-of-distribution) are associated with smaller “confidence” scores such as softmax probability (Hendrycks and Gimpel 2017; Northcutt, Jiang, and Chuang 2021), maximum logit (Hendrycks et al. 2022; Wei et al. 2022), $\ell^{2}$-distance (Sun et al. 2022), energy score (Liu et al. 2020b), and cosine similarity (Ngoc-Hieu et al. 2023). Recently, LLM has shown a tendency to produce a lower (higher) confidence in its incorrect (correct) answers in multiple-choice Q&A (Plaut, Nguyen, and Trinh 2024). Building on previous works, we hypothesized that the logit of generated tokens by RMU models exhibit randomness. As seen by a deep neural network, such randomization signifies low confidence in the logit, resulting in nonsensical or incorrect responses. To evaluate our hypothesis, we conducted a theoretical analysis of the logits of generated tokens produced by RMU models. To facilitate subsequent analysis, we make the following definition and assumption.
@@
(Unlearned model & logit of tokens on unlearned model). Let $f^{k,l}=g^{(k)}\circ h^{(l)}$ be the transformation from layer $l$ to layer $k$ of network $f$, for any two layers $k\geq l$; $l\in[1...L]$, $k\in[l+1...L]$. We define the unlearned model $f^{\textnormal{unlearn}}=\bm{W}(g^{(L)}\circ h^{(l),\textnormal{steered}}$), where $h^{(l),\textnormal{steered}}(x_{F})$ is the steered representation of forget input $x_{F}$ at layer $l$ and $\bm{W}$ is the unembedding matrix. Given a prompt input $x_{F,1:n}$. For a next token $x_{n+1}$, the logit of $x_{n+1}$ obtained from unlearned model $f^{\textnormal{unlearn}}$ is defined as:
$f^{\textnormal{unlearn}}(x_{n+1}|x_{F,1:n})$ $=\bm{W}(g^{(L)}\circ h^{(l),\textnormal{steered}})(x_{n+1}|x_{F,1:n})$
$=\bm{W}g^{(L)}(h^{(l),\textnormal{steered}}(x_{n+1}|x_{F,1:n}))$ (2)
@@
A well-unlearned model shifts the representation $h^{(l),\textnormal{steered}}(x_{F})$ of a forget-sample $x_{F}$ at layer $l$ to a scaled random vector $c\bm{u}$. More concretely,
$h^{(l),\textnormal{steered}}(x_{F})=c\bm{u}+\bm{\epsilon},$ (3)
where $\bm{\epsilon}$ is a small error. Without losing generality, we assume that $\bm{\epsilon}$ is sampled from Normal distribution $\mathcal{N}(\bm{0},\eta\bm{I})$, where $\eta\bm{I}$ is the covariance matrix, $\eta\in\mathbb{R}$.
@@
If Assumption 1 holds, by Definition 1, the logit of token $x_{n+1}$ generated by unlearned model $f^{\textnormal{unlearn}}$ given as $f^{\textnormal{unlearn}}(x_{n+1}|x_{F,1:n})$ follows the Normal distribution $\mathcal{N}\left(\bm{W}g^{(L)}(\bm{z}),\eta\bm{W}\nabla_{\bm{z}}g^{(L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(L)}(\bm{z})\bm{W}^{\top}\right),$ where $\bm{z}=c\bm{u}$.
@@
Given Assumption 1, we have:
$h^{(l),\textnormal{steered}}(x_{n+1}|x_{F,1:n})=c\bm{u}+\bm{\epsilon},$ (4)
We denote $\bm{z}=c\bm{u}$. Substituting Eqn. 3 into Eqn. 2, we get:
${f^{\text{unlearn}}}(x_{n+1}|x_{F,1:n})=\bm{W}g^{(L)}(\bm{z}+\bm{\epsilon})$ (5)
Since $\bm{\epsilon}$ is small, we approximate the function $g^{(L)}(\bm{z}+\bm{\epsilon})$ by its first-order derivative:
$f^{\text{unlearn}}(x_{n+1}|x_{F,1:n})$ $=\bm{W}g^{(L)}(\bm{z}+\bm{\epsilon})$
$\approx\bm{W}g^{(L)}(\bm{z})+\bm{W}\nabla_{\bm{z}}g^{(L)}(\bm{z})^{\top}\bm{\epsilon}$ (6)
@@
$f^{\textnormal{unlearn}}$ $(x_{n+1}|x_{F,1:n})$
$\sim\mathcal{N}(\bm{W}g^{(L)}(\bm{z}),\eta\bm{W}\nabla_{z}g^{(L)}(\bm{z})^{\top}\nabla_{z}g^{(L)}(\bm{z})\bm{W}^{\top})$ (7)
Since $\bm{u}\sim U(0,1)$, then $c\bm{u}\sim U(0,c)$. By the linearity property of expectation and definition of variance, we have: $\mathbb{E}(\bm{z})=\mathbb{E}(c\bm{u})=\frac{c}{2}$; $\textnormal{Var}(\bm{z})=\textnormal{Var}(c\bm{u})=c^{2}\textnormal{Var}(\bm{u})=\frac{c^{2}}{12}$
Proposition 1 suggests that he variance of $f^{\textnormal{unlearn}}(x_{n+1}|x_{F,1:n})$ is controlled by (i) $\eta$: a scalar variance and (ii) $\bm{W}\nabla_{\bm{z}}g^{(L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(L)}(\bm{z})\bm{W}^{\top}$: the product of $\bm{W}\nabla_{\bm{z}}g^{(L)}(\bm{z})^{\top}$ and $\nabla_{\bm{z}}g^{(L)}(\bm{z})\bm{W}^{\top}$. If $f^{\textnormal{unlearn}}(x_{n+1}|x_{F,1:n})$ has high variance, the logit values are more random. Since $\bm{\epsilon}$ presents a small error, then $\bm{\epsilon}$ is vary for different input $x_{F}$. This variation makes it difficult to control the variance of the logit by $\eta$. The main effect depend on $\bm{W}\nabla_{\bm{z}}g^{(L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(L)}(\bm{z})\bm{W}^{\top}$. While the unembedding matrix $\bm{W}$ is unchanged after unlearning, the product $\nabla_{\bm{z}}g^{(L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(L)}(\bm{z})$ is vary depends on the specific characteristics of sub-networks $g^{(L)}$ and input $\bm{z}=c\bm{u}$. Unfortunately, $g^{(L)}$ is a composition of transformer layers, which is highly nonlinear, making it difficult to have a complete analysis. The variance of $\bm{z}$ is derived as $\textnormal{Var}(\bm{z})=\frac{c^{2}}{12}$. When $c$ gets larger, the variance of $\bm{z}$ is higher. This could increase the variability of $g^{(L)}(\bm{z})$ and the gradient $\nabla_{\bm{z}}g^{(L)}(\bm{z})$. A larger $c$ could introduces more randomness to the logit. We conduct an empirical analysis to understand the confidence of generated tokens by RMU models in Section 4.1.

### v2

@@ Abstract
Representation Misdirection for Unlearning (RMU), which steers model representation in the intermediate layer to a target random representation, is an effective method for large language model (LLM) unlearning. Despite its high performance, the underlying cause and explanation remain underexplored. In this paper, we theoretically demonstrate that steering forget representations in the intermediate layer reduces token confidence, causing LLMs to generate wrong or nonsense responses. We investigate how the coefficient influences the alignment of forget-sample representations with the random direction and hint at the optimal coefficient values for effective unlearning across different network layers. We show that RMU unlearned models are robust against adversarial jailbreak attacks. Furthermore, our empirical analysis shows that RMU is less effective when applied to the middle and later layers in LLMs. To resolve this drawback, we propose Adaptive RMU—a simple yet effective alternative method that makes unlearning effective with most layers. Extensive experiments demonstrate that Adaptive RMU significantly improves the unlearning performance compared to prior art while incurring no additional computational cost.
@@ 1 Introduction
LLMs achieved remarkable performance through pre-training on large amounts of internet texts and rigorous alignment processes for safety enhancement. Despite the immense effort in safety research, LLMs are still vulnerable to adversarial jailbreak attacks and can exhibit unwanted behaviors (Shah et al. 2023; Zou et al. 2023b; Jones et al. 2023; Yuan et al. 2024; Wei, Haghtalab, and Steinhardt 2024).
Machine Unlearning (Cao and Yang 2015; Bourtoule et al. 2021; Nguyen et al. 2022; Xu et al. 2023; Liu et al. 2024c) has emerged as a promising method for mitigating unforeseen risks in LLMs before deployment. Li et al. (2024b) introduced Representation Misdirection for Unlearning (RMU)—an unlearning method that steers the representations of forget-samples (i.e. samples that the model should forget) toward a random representation while keeping the representations of retain-samples (i.e. samples that the model should remember) unchanged. RMU significantly degrades models’ accuracy on forget-tasks, while only slightly affecting the performance on retain-tasks and demonstrates stronger robustness against adversarial jailbreak attacks. However, the reason for RMU’s effectiveness is not well understood, hindering the development of better unlearning algorithms. In this paper, we make the following contributions:
@@
  • We investigate the connection between RMU and adversarial robustness. We demonstrate that RMU impedes the adversary’s ability to determine optimal updates for generating adversarial samples, thus improving the adversarial robustness of the unlearned model.
@@ 2 Background and related work
A natural unlearning approach is leave-some-out retraining: retraining the model from scratch without the forget samples. However, this method becomes more computationally expensive as the size of datasets and modern deep networks grows. Existing works focus on approximating unlearning (Warnecke et al. 2021; Izzo et al. 2021; Sekhari et al. 2021; Isonuma and Titov 2024) using influence function (Koh and Liang 2017; Grosse et al. 2023), gradient ascent (Thudi et al. 2022), second-order approximation (Jia et al. 2024), negative preference optimization (Zhang et al. 2024b), and embedding corrupted (Liu et al. 2024a). Other views on the landscape of machine unlearning include: unlearning in text classification (Ma et al. 2022), image classification and recognition (Ginart et al. 2019; Golatkar, Achille, and Soatto 2020; Fan et al. 2024; Choi and Na 2023; Cha et al. 2024), image-to-image generative models (Li et al. 2024a), diffusion models (Gandikota et al. 2023; Zhang et al. 2024a; Kumari et al. 2023; Bui et al. 2024), multimodal unlearning (Cheng and Amiri 2023), federated unlearning (Romandini et al. 2024; Wang et al. 2022; Che et al. 2023; Halimi et al. 2022; Jeong, Ma, and Houmansadr 2024), graph unlearning (Chen et al. 2022; Chien, Pan, and Milenkovic 2023; Wu et al. 2023a; Cheng et al. 2023; Dukler et al. 2023; Zhu, Li, and Hu 2023; Li et al. 2024c; Tan et al. 2024), recommender systems (Zhang et al. 2023; Chen et al. 2024; Li et al. 2023; Wang et al. 2025), certified minimax unlearning (Liu et al. 2024b), targeted types of unlearning information (Cooper et al. 2024), and evaluation on unlearning (Lynch et al. 2024; Hayes et al. 2024; Shi et al. 2024a, b).
@@
Due to the large size of the parameters and training data, LLM poses a new challenge to unlearning. Recent studies in LLM unlearning mainly focus on task or context-specific settings such as unlearning copyrighted material from the Harry Potter series (Eldan and Russinovich 2023), in-context unlearning (Pawelczyk, Neel, and Lakkaraju 2024), fictitious unlearning (Maini et al. 2024), specific harmful input-output (Yao, Xu, and Liu 2023; Liu et al. 2024d), sensitive and private information (Jang et al. 2023; Wu et al. 2023b; Patil, Hase, and Bansal 2024), gender bias (Belrose et al. 2023) or concepts (Hong et al. 2024; Bui et al. 2024). More recently, Li et al. (2024b) consider unlearning an entire distribution of hazardous knowledge given limited samples.
@@
Let $\mathcal{D}_{\text{forget}}$ and $\mathcal{D}_{\text{retain}}$ be the forget and retain sets, respectively. Let $f_{\theta}:\mathbb{R}^{n\times d}\mapsto\mathbb{R}^{n\times|V|}$ be an autoregressive LLM parameterized by $\theta$ that maps a prompt input $x_{1:n}$ consisting of $n$ tokens $\{x_{1},x_{2},...,x_{n}\}$ to an output of probability distributions over the vocabulary $V$. We denote $h_{\theta}^{(l)}(x)$ the averaged hidden states of all tokens in $x_{1:n}$ obtained from the $l$-th layer of $f_{\theta}$. For simplicity, throughout this paper, we use $h^{(l)}(x)$ to present $h_{\theta}^{(l)}(x)$. For operators, we denote $\circ$ as the decomposition operator, and |||||$\cdot$||| | ⋅ | | is the Euclidean norm. Our goal is to unlearn the undesired harmful knowledge $\mathcal{D}_{\textnormal{forget}}$ from $f_{\theta}$ while retaining general knowledge $\mathcal{D}_{\textnormal{retain}}$. Unlearned models should be robust to knowledge recovery attacks that attempt to recover harmful knowledge from the model.
Representation Misdirection for Unlearning
(RMU; Li et al. (2024b)) is a fine-tuning based unlearning method inspired by representation engineering (Zou et al. 2023a) that steers the model’s representation of forget samples $x_{F}\in\mathcal{D}_{\text{forget}}$ to a random vector and regularizes the model representation of retain samples $x_{R}\in\mathcal{D}_{\text{retain}}$ back to the original model representation, by optimizing the MSE loss:
$\mathcal{L}$ $=\mathbb{E}_{x_{F}\in\mathcal{D}_{\text{forget}}}||h_{\theta^{\text{unlearn}}}^{(l)}(x_{F})-c\bm{u}||_{2}^{2}$
$+\alpha\mathbb{E}_{x_{R}\in\mathcal{D}_{\text{retain}}}||h_{\theta^{\text{unlearn}}}^{(l)}(x_{R})-h_{\theta^{\text{frozen}}}^{(l)}(x_{R})||_{2}^{2},$ (1)
where $\theta^{\text{unlearn}}$ and $\theta^{\text{frozen}}$ are parameters of the update model and frozen model respectively, $\bm{u}$ is a fixed random unit vector where each element is sampled from Uniform distribution $U(0,1)$, $c\in\mathbb{R}$ is a fixed scaling coefficient and $\alpha\in\mathbb{R}$ is a retain weight. RMU updates $\theta^{\text{unlearn}}$ toward the direction of the gradient of the loss $\mathcal{L}$ using gradient descent.
@@ 3.1 The confidence of tokens generated by RMU models
In general, samples from the shifted distribution (such as wrong label or out-of-distribution) are associated with smaller “confidence” scores such as softmax probability (Hendrycks and Gimpel 2017; Northcutt, Jiang, and Chuang 2021), maximum logit (Hendrycks et al. 2022; Wei et al. 2022), $\ell^{2}$-distance (Sun et al. 2022), energy score (Liu et al. 2020), and cosine similarity (Ngoc-Hieu et al. 2023). Recently, LLM has shown a tendency to produce a lower (higher) confidence in its incorrect (correct) answers in multiple-choice Q&A (Plaut, Nguyen, and Trinh 2024). Building on previous works, we hypothesized that the logit of generated tokens by RMU models exhibit randomness. As seen by a deep network, such randomization signifies low confidence in the logit, resulting in nonsensical or incorrect responses. To validate the hypothesis, we conducted an analysis of the logits of generated tokens produced by RMU models. To facilitate subsequent analysis, we make the following definition and assumption.
@@
(Unlearned model & logit of forget-tokens on unlearned model). Let $f^{(l:k)}=g^{(l:k)}\circ h^{(l)}$, where $g^{(l:k)}$ be the transformation from layer $l$ to layer $k$ of network $f$, for any two layers $k>l$; $l\in[1...L]$. We define the unlearned model $f^{\textnormal{unlearn}}=\bm{W}(f^{(l:L),\textnormal{steered}})=\bm{W}(g^{(l:L)}\circ h^{(l),\textnormal{steered}})$, $h^{(l),\textnormal{steered}}$ is the steered representation of the given input at layer $l$ and $\bm{W}$ is the unembedding matrix which maps output hidden states back to the vocabulary space. Given a forget input $x_{F,1:n}$, the logit of the next token $x_{F,n+1}$ obtained from unlearned model $f^{\text{unlearn}}$ is defined as:
$f^{\text{unlearn}}($ $x_{F,n+1}|x_{F,1:n})=\bm{W}f^{(l:L),\text{steered}}(x_{F,n+1}|x_{F,1:n})$
$=\bm{W}(g^{(l:L)}\circ h^{(l),\textnormal{steered}})(x_{F,n+1}|x_{F,1:n})$
$=\bm{W}g^{(l:L)}(h^{(l),\textnormal{steered}}(x_{F,n+1}|x_{F,1:n}))$ (2)
@@
A well-unlearned model shifts the representation of all tokens in a forget-sample $x_{F,1:n}$ at layer $l$ to a scaled random vector $c\bm{u}$. More concretely,
$h^{(l),\textnormal{steered}}(x_{F,i})=c\bm{u}+\bm{\epsilon},$ (3)
where $x_{F,i}$ is the $i$-th token in $x_{F}$, $\bm{\epsilon}$ is a small error. Without losing generality, we assume that $\bm{\epsilon}$ is sampled from Normal distribution $\mathcal{N}(\bm{0},\eta\bm{I})$, where $\eta\bm{I}$ is the covariance matrix, $\eta\in\mathbb{R}$.
@@
If Assumption 1 holds, by Definition 1, the logit value of forget token $x_{F,n+1}$ generated by unlearned model $f^{\textnormal{unlearn}}$ given as $f^{\textnormal{unlearn}}(x_{F,n+1}|x_{F,1:n})$ follows the Normal distribution $\mathcal{N}\left(\bm{W}g^{(l:L)}(\bm{z}),\eta\bm{W}\nabla_{\bm{z}}g^{(l:L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(l:L)}(\bm{z})\bm{W}^{\top}\right),$ where $\bm{z}=c\bm{u}$.
@@
Assumption 1 implies that in a well-unlearned model, token $x_{F,n+1}$ is independent of the previous tokens, thus we have:
$h^{(l),\textnormal{steered}}(x_{n+1}|x_{F,1:n})\approx h^{(l),\text{steered}}(x_{F,n+1})=c\bm{u}+\bm{\epsilon}$ (4)
Denote $\bm{z}=c\bm{u}$. Substituting Eqn. 4 into Eqn. 2, we get:
${f^{\text{unlearn}}}(x_{F,n+1}|x_{F,1:n})\approx\bm{W}g^{(l:L)}(\bm{z}+\bm{\epsilon})$ (5)
Since $\bm{\epsilon}$ is small, we approximate the function $g^{(l:L)}(\bm{z}+\bm{\epsilon})$ by its first-order derivative:
$f^{\text{unlearn}}(x_{F,n+1}|x_{F1:n})\approx\bm{W}(g^{(l:L)}(\bm{z})+\nabla_{\bm{z}}g^{(l:L)}(\bm{z})^{\top}\bm{\epsilon})$ (6)
@@
$f^{\textnormal{unlearn}}(x_{F,n+1}|x_{F,1:n})$
$\sim\mathcal{N}\left(\bm{W}g^{(l:L)}(\bm{z}),\eta\bm{W}\nabla_{z}g^{(l:L)}(\bm{z})^{\top}\nabla_{z}g^{(l:L)}(\bm{z})\bm{W}^{\top}\right)$ (7)
Since $\bm{u}\sim U(0,1)$, then $\bm{z}\sim U(0,c)$. By definition of variance, we have: $\textnormal{Var}(\bm{z})=\textnormal{Var}(c\bm{u})=c^{2}\textnormal{Var}(\bm{u})$. ∎
Proposition 1 suggests that the variance of $f^{\text{unlearn}}(x_{F,n+1}|x_{F,1:n})$ is controlled by (i) $\eta$: a scalar variance and (ii) $\bm{W}\nabla_{\bm{z}}g^{(l:L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(l:L)}(\bm{z})\bm{W}^{\top}$: the product of $\bm{W}\nabla_{\bm{z}}g^{(L)}(\bm{z})^{\top}$ and $\nabla_{\bm{z}}g^{(L)}(\bm{z})\bm{W}^{\top}$. If $f^{\text{unlearn}}(x_{F,n+1}|x_{F,1:n})$ has high variance, the logit values are more random. Since $\bm{\epsilon}$ presents a small error, then $\bm{\epsilon}$ varies for different inputs $x_{F}$. This variation makes it difficult to control the variance of the logit by $\eta$. The main effect depend on $\bm{W}\nabla_{\bm{z}}g^{(l:L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(l:L)}(\bm{z})\bm{W}^{\top}$. While the unembedding matrix $\bm{W}$ is unchanged after unlearning, the product $\nabla_{\bm{z}}g^{(l:L)}(\bm{z})^{\top}\nabla_{\bm{z}}g^{(l:L)}(\bm{z})$ varies depending on the specific characteristics of sub-networks $g^{(l:L)}$ and input $\bm{z}=c\bm{u}$. Unfortunately, $g^{(l:L)}$ is a composition of transformer layers, which is highly nonlinear, making it difficult to have a complete analysis. The variance of $\bm{z}$, derived as $\text{Var}(\bm{z})=c^{2}\text{Var}(\bm{u})$, is proportional to $c$; i.e. when $c$ gets larger, the variance of $\bm{z}$ is higher. This could increase the variability of $g^{(l:L)}(\bm{z})$ and the gradient $\nabla_{\bm{z}}g^{(l:L)}(\bm{z})$. A larger $c$ could introduces more randomness to the logit. We conduct an empirical analysis to understand the confidence of generated tokens by RMU models in Section 4.1.